OpenStudy (turingtest):

Not Fool's problem of the day: Prove\[\int_{0}^{\pi}\frac{x\sin x}{1+\cos^2x}dx=\frac{\pi^2}{4}\]

OpenStudy (turingtest):

This has already been solved, so you could probably find the answer in my profile, but I wanna see a different approach.

OpenStudy (anonymous):

let's check out your profile then ... LOL

OpenStudy (experimentx):

intergration by parts and subs cosx=y ???

OpenStudy (turingtest):

no, tha would be cheating :( use your own method

OpenStudy (anonymous):

ehh.. don't worry, I have no idea about integrals. I can't do it. No worries ;)

OpenStudy (turingtest):

integration by parts is a good start @experimentX

OpenStudy (experimentx):

i don't think i will be posting ... if i'ever post ... i will post one or two steps before solution. I am terrible at latex.

OpenStudy (anonymous):

Lol I love the name! :D

OpenStudy (turingtest):

@FoolForMath sorry, didn't mean to steal your moment :P a coincidence I assure you!

OpenStudy (anonymous):

Lolz sure mate :)

OpenStudy (ash2326):

We have \[I= \int_{0} ^{\pi} \frac{x \times \sin x}{1+\cos^2 x} dx..................(1)\] We know differential integral property \[I= \int_a ^b f(x) dx=\int_a ^b (a+b-x) dx\] so \[I= \int_{0} ^{\pi} \frac{(\pi-x) \times \sin(\pi- x)}{1+\cos^2 (\pi-x)} dx\] we get now \[I= \int_{0} ^{\pi} \frac{(\pi-x) \times \sin x}{1+\cos^2 x} dx..................(2)\] Add 1 and 2 , we get \[2I= \int _{0} ^{\pi} \frac{(\pi) \times \sin x}{1+\cos^2 x} dx\] now substitute cos x= u and solve !!!

OpenStudy (experimentx):

a way lot better method than i would have imagined ... everybody has one specialty. i guess for @ffm it smelling numbers.

OpenStudy (anonymous):

Lol, I love numbers!

OpenStudy (anonymous):

Someone stop this @Fweaa guy !!

OpenStudy (ash2326):

@Fweaa please don't spam , post your question only once and not in someone else's question!!!!

OpenStudy (anonymous):

Turing bring in the hammer :P

OpenStudy (turingtest):

awesome, quite different than Zarkon's approach, which I will now post and I'l get this Fw whatyever guy ;)

OpenStudy (anonymous):

@ash2326: Amazing! funny thing I understand your approach :P

OpenStudy (ash2326):

Thanks @FFM

OpenStudy (anonymous):

The silly thing is my friends who are not great in mathematics finds integration so easy but I always find it difficult after high school :(

OpenStudy (turingtest):

here is Zarkons solution: http://openstudy.com/users/turingtest#/updates/4f047463e4b075b56651c384 @ash2326 thanks for another approach :) I do see similarities though

OpenStudy (anonymous):

Turing: What's up with the Fweaa??!!

OpenStudy (experimentx):

thats what i would have done. --> the classic way.

OpenStudy (anonymous):

Everybody would had solved this except me :P

OpenStudy (ash2326):

@TuringTest I'll continue the solution \[2I=\pi \int_0 ^{\pi} \frac{\sin x}{1+\cos ^2 x} dx\] u= cos x so du =-dx sin x x=0 u=1 x=pi u=-1 \[2I=\pi \int _{1}^{-1} \frac{- du}{1+u^2}\] so we have now \[2I=\pi \int _{-1} ^{1} \frac{du}{1+u^2}\] we know for an even function \[\int_{-a} ^{a} f(x) dx= 2\int_0 ^{a} f(x) dx\] so we get \[2I=2\pi \int_0 ^{1} \frac{du}{1+u^2} du\] or \[I= \pi \times (\tan^{-1} 1)= \frac{\pi^2}{4}\]

OpenStudy (experimentx):

@ffm you didn't try at all, you would definitely, but the way ash2326 done is genius ... i'll remember it.

OpenStudy (turingtest):

sorry, took Fwee is the first I've suspended, so I had to figure out exactly how it works @ash2326 awesome! quite different from the other way :)

OpenStudy (anonymous):

Lol, Mate I was good at it since high school ;) I lost interest after I met number theory and combinatoric in clg. But may be that's an lame excuse :(

OpenStudy (experimentx):

you are in different league. I am still struggling with calculus.

OpenStudy (ash2326):

Thanks @TuringTest

OpenStudy (turingtest):

I just love integrals. This was a fun one. PS: @experimentX I don't think you're tagging FoolForMath by writing "@ffm"

OpenStudy (anonymous):

Hey guys! I just opened my book and I saw this problem, exactly same approach as ash!

OpenStudy (experimentx):

he's here ... besides @ffm is more convenient than using whole name.

OpenStudy (ash2326):

Yeah, this approach is taught everywhere in India:D

OpenStudy (anonymous):

Sure @experiment!! anything fool, ffm whatever you want :P

OpenStudy (anonymous):

Okay I think I got a good problem for you guys :P I would post it in a new thread ;) Ash you are not supposed to use our indigenous trick :P

OpenStudy (ash2326):

Yeah:D:D I shouldn't have:P

OpenStudy (callisto):

Accidentally, I discovered that it was one of my mock exam question.. and my teacher did it like that.